OpenStudy (anonymous):

Some one please helpp The table below shows two equations: Equation 1 |3x - 1| + 7 = 2 Equation 2 |2x + 1| + 4 = 3 Which statement is true about the solution to the two equations?

OpenStudy (texaschic101):

| 3x - 1 | + 7 = 2 -- subtract 7 from both sides | 3x - 1 | = 2 - 7 | 3x - 1 | = -5 3x - 1 = -5 3x = -5 + 1 3x = -4 x = -4/3 and -(3x - 1) = -5 -3x + 1 = -5 -3x = -5 -1 -3x = -6 x = 2 solutions are : x = -4/3 and x = 2 ======================== | 2x + 1 | + 4 = 3 -- subtract 4 from both sides | 2x + 1 | = 3 - 4 | 2x + 1 | = -1 2x + 1 = -1 2x = -1 - 1 2x = -2 x = -1 -(2x + 1) = -1 -2x - 1 = -1 -2x = -1 + 1 -2x = 0 x = 0 solutions are : x = -1 and x = 0 =========================== now what are your answer choices ?

OpenStudy (true_alpha_wolf):

DIY